Problem1 – Process costing


Timing Tech, a manufacturing company, has two departments: Production department and Finishing department. Goods completed in the Production department are immediately sent to the Finishing department. The following data pertains to the Production department.


Physical Direct Material Direct Labor

Units Cost Cost

Beg. WIP inventory (Jan. 1 2014) 200,000 units $ 200,000 $ 315,000

Units Started during the year 1,000,000 units

Costs incurred during the year $1,300,000 $1,995,000

End. WIP inventory (Dec. 31 2014) 300,000 units


  • Beg. WIP completion rate: Direct Material – 100%, Conversion cost – 50%

  • End. WIP completion rate: Direct Material – 100%, Conversion cost – 80%

  • The company uses normal costing for allocation of manufacturing overhead. The pre-determined allocation rate in 2013 and 2014 is 60% of direct labor cost.

Required) Answer the following questions using a) FIFO method and b) Weighted Average Method.

  1. FIFO


  1. What is the number of physical units for the following accounts:


Goods Completed ______________

WIP End. Balance ___________________


  1. What is the cost per equivalent unit?


Direct Material $_______________

Conversion $_______________



  1. Compute the cost assigned to goods completed/transferred-out to testing department and the cost assigned to WIP ending balance for the Production department.


Transferred-out $_________________

WIP ending balance $________________


Problem 1 – Continued


  1. Weighted average method


  1. What is the number of physical units for the following accounts:


Goods Completed ______________

WIP End. Balance ___________________


  1. What is the cost per equivalent unit?


Direct Material $_______________

Conversion $_______________



  1. Compute the cost assigned to goods completed/transferred-out to testing department and the cost assigned to WIP ending balance for the Production department.


Transferred-out $_________________

WIP ending balance $________________
Answer #1

1 Statement showing physical units

Units to account for
Beginning work in process on Jan 1, 2014 200000
Started during the period 1000000
Total physical units to account for 1200000
Units accounted for :
Goods completed 900000 (1200000-300000)
WIP ending balance 300000

2 . Statement showing equivalent units of production :-

Physical units equivalent units Equivalent units
Units to account for : Direct material direct labor
Beginning WIP 200000
Started during the period 1000000
Units to account for 1200000
Beginning WIP completed 200000 0 100000 (200000x50%)
Started and transferred 700000 700000 (700000x100%) 700000 (700000x100%)
Ending WIP 300000 300000 (300000x100%) 240000 (300000x80%)
Total physical Units accounted for 1200000
Equivalent units 1000000 1040000

Cost per equivalent units = cost incurred during the period (2014) / equivalent units

Direct material = $1300000 / 1000000 = $1.3

Direct labor = $1995000/1040000 = $1.92

3. Cost assigned to completed and transferred out = 700000x1.3+800000x1.92 = 910000+1534615 = $2444615.

Cost assigned to WIP ending balance = 300000x1.3+240000x1.92 = 390000+460385 = $850385.

Weighted average

1. Statement showing physical units :-

Units to account for :
Beginning WIP 200000
Started during the year 1000000
Total physical units to account for 1200000
Units accounted for :-
Goods completed

900000

(1200000-300000)

Ending WIP inventory 300000

2. Statement showing equivalent units produced :-

Equivalent units equivalent units
Units to account for :- Physical units direct material direct labor
Beginning WIP 200000
Started during the period 1000000
Total physical units to be accounted for 1200000
Completed and transferred 900000 900000 (900000x100%) 900000 (900000x100%)
Ending WIP 300000 300000 (300000x100%) 240000 (300000x80%)
Total physical units accounted for 1200000
Equivalent units 1200000 1140000

Equivalent cost = beginning cost + cost incurred during the period / equivalent units

= Direct material = 200000+1300000/1200000 = $1.25

Direct labor = 315000+1995000/1140000 = $2.03

3. Cost assigned to gto giods completed and transferred out = 900000x1.25+900000x2.03 = 1125000+1823684 = $2948684 or

Cost assigned to ending WIP = 300000x1.25+240000x2.03 = 375000+486316 = $861316.